A 15.2 kg block is dragged over a rough, horizontal surface by a constant force of 133 N acting at an angle of 34.2 ◦ above the horizontal. The block is displaced 50.6 m, and the coefficient of kinetic friction is 0.176. Find the magnitude of the work done by the force of friction. 


1
Expert's answer
2021-11-19T10:50:31-0500

By the definition of the work done, we have:


"W_{fr}=F_{fr}d."

The friction force can be found as follows:


"F_{fr}=\mu_kN=\mu_k(mg-F_{appl}sin\theta)."

Substituting the friction force into the previous equation, we get:


"W_{fr}=\mu_kd(mg-F_{appl}sin\theta),""W_{fr}=0.176\times50.6\ m\times(15.2\ kg\times9.8\ \dfrac{m}{s^2}-133\ N\times sin34.2^{\circ}),""W_{fr}=661\ J."